OpenStudy (anonymous):

divide: 4x^4-15X^3-5x^2-12x+1 / 3x

OpenStudy (cruffo):

Do you mean: \[\large \frac{4x^4-15x^3-5x^2-12x+1}{3x}\]

OpenStudy (anonymous):

yea

OpenStudy (cruffo):

You don't need long division with this one. If there is one one term in the bottom, split the top up across + and -: \[ \large \frac{4x^4}{3x} - \frac{15x^3}{3x} - \frac{5x^2}{3x} - \frac{12x}{3x} + \frac{1}{3x}\] Then simplify the individual fractions.

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

i got -19/3 x^8

OpenStudy (cruffo):

hmmm.... not sure what your doing... I got \[\large \frac{4}{3}x^3 - 5x^2 - \frac{5}{3}x - 4 + \frac{1}{3x}\]

OpenStudy (cruffo):

Sorry "your" should be "you're" :)

OpenStudy (cruffo):

I think you may be trying to "add like terms", but they are not "like terms"

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

yea that's what i am doing

OpenStudy (anonymous):

can we do a couple more

OpenStudy (cruffo):

I will be logging off in a few minutes. Just keep posting your questions.